Grade 11Physical Chemistry

5 moles of Na2S2O3 loses 40 moles of electron . Determine new oxidation no. of sulphur

Aryan Tripathi
7 Years agoGrade 11

1 Answer

Pooja
7 Years ago
Here the combined oxidation no of both S atoms would be
2(Na)+2(S)+3(O)=0
2(1)+2(S)+3(-2)=0
2(S)=+4
 
1 mole has +4 
5 mole will have ON= +20
 
5S2O3-2  = 10SO4-2  + 40e-     
+20                                -40
It is clear from the above rection that the ON of 10 mole product will be 20 - (-40) = +60
So ON of one mole will be +6
 
S